A house which is valued at ksh2600,000 appreciated at rate of 12% per year.



I. What will be its value after two and half years?



Ii. After how long will its value be ksh 4200,000





1
Expert's answer
2021-12-06T12:33:12-0500

I. Value of house after two and half years

"FV=PV(1+r)^{n}"


"FV=2600000\times(1+0.12)^{2.5}"

"FV=3,451,583.66"


Ii. Time period to reach 4200,000

"FV=PV(1+r)^{n}"


"4200000=2600000\times(1+0.12)^{n}"


"(1+0.12)^{n}=\frac{4200,000}{2600,000}"


"(1.12)^{n}=1.62"


"n\times log(1.12)=log(1.62)"


"n=\frac{log(1.62)}{log(1.12)}"


"n=\frac{0.2095}{0.0492}=4.26"


This means the value of house will be ksh 4200,000 after 4.26 years.